# Global Extraction of the $\rm{}^{12}C$ Nuclear Electromagnetic Response Functions (${\cal R}_{L}$ and ${\cal R}_{T}$) and Comparisons to Nuclear Theory and Neutrino/Electron Monte Carlo Generators

### v2

@@ I Introduction
Electron scattering cross sections on nuclear targets are completely described by longitudinal (${\cal R}_{L}$) and transverse (${\cal R}_{T}$) nuclear electromagnetic response functions. Here ${\cal R}_{L}$ and ${\cal R}_{T}$ are functions of the energy transfer $\nu$ (or excitation energy $E_{x}$) and the square of the 4-momentum transfer $Q^{2}$ (or alternatively the 3-momentum transfer $\bf q$). Recent theoretical  Mihaila and Heisenberg (2000); Lovato et al. (2016); Cloët et al. (2016); Franco-Munoz et al. (2022, 2023); Sobczyk et al. (2020); Carlson et al. (2015, 2002) of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ can be tested by comparing the predictions to experimental data. We perform extraction of both response functions as functions of (${\bf q},\nu$), as well as ($Q^{2},\nu)$ from an analysis of all available inclusive cross section data for $\rm{}^{12}C$, and compare them to theoretical predictions.
@@
In this paper we focus on the nuclear excitation, quasielastic (QE) and two nucleon final state (2p2h) regions and compare the extracted $\rm{}^{12}C$ ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ to nuclear theory predictions of the “Green’s Function Monte Carlo” (GFMC) Lovato et al. (2016), the “Energy Dependent-Relativistic Mean Field” (ED-RMF) Franco-Munoz et al. (2022, 2023), the “Short Time Approximation Quantum Monte Carlo” (STA-QMC) Andreoli et al. (2024), the “Correlated Fermi Gas” (CFG) Bhattacharya et al. (2024), the NuWro theoretical approach NuW (2024), the achilles Isaacson et al. (2023) theoretical approach, and an improved superscaling model (SuSAv2)Gonzaléz-Jiménez et al. (2014); Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023). We have not done comparisons to the gibuu theoretical approach because nuclear response function predictions by gibuu are not readily available.
The theoretical approaches of NuWro and achilles  have been implemented in the corresponding neutrino MC generators. Recently, ED-RMF has been implementedMcKean (2024) in an update of the neut neutrino MC generator. An earlier version of SuSAv2 (QE+2p2h) has been implemented in genie, SuSAv2-inelastic is being implemented in genie, and STA-QMC for $\rm{}^{4}He$ only has been implemented in genie. In this paper we focus on testing theoretical predictions. The testing of the implementation of the theoretical approaches into the genie neutrino MC generator (run in electron scattering mode) will be presented in a future communication. Similar investigation of electron scattering data on $\rm{}^{40}Ca$ and $\rm{}^{56}Fe$ (as well as hydrogen and deuterium data) are currently under way.
@@
In this communication we report on the extraction of ${\cal R}_{L}$ and ${\cal R}_{T}$ for ${\rm{}^{12}C}$ at 18 distinct values of both q and $Q^{2}$ by including all available electron scattering cross section measurements on carbon. We note that a significant number of cross section measurements were not available in tabular form and required digitization of the cross sections from figures in the publications. As summarized in Table 1, the response functions are extracted for a large range of energy transfer $\nu$ at 18 fixed values of $\bf q$ in the range $3.750.1<{\bf q}<$ 3.75  GeV and at 18 fixed values of $Q^{2}$ in the range $3.50<Q^{2}<$ 3.5 GeV2. The range of $\nu$ spans the nuclear excitation, QE, resonance and inelastic continuum. The complete data set consists of about 10,000 ${\rm{}^{12}C}$ differential cross section measurements points as well as photo-absorption data ($Q^{2}=0$). We also include high precision cross section measurements from Jefferson Lab Hall C experiment E-04-001 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019).
@@
In the nuclear excitation region (excitation energy less than 50 MeV) we extract ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from fits to ${\rm{}^{12}C}$ nuclear excitation form factors by Bodek and Christy Bodek and Christy (2023). In addition, for $4016<E_{x}<$ 40 MeV, values of ${\cal R}_{L}({\bf q},E_{x})$ and ${\cal R}_{T}({\bf q},E_{x})$ are available from the analysis of Yamaguchi:71 Yamaguchi et al. (1971) for five values of $\bf q$ (0.148, 0.167, 0.205, 0.240 and 0.307 GeV). Comparisons of the fit to the Yamaguchi response functions and cross sections are shown in Figures 1 and 2, respectively.
@@ I.1 Overview of this analysis
The Christy-Bodek universal fit to all available cross section measurements for ${\rm{}^{12}C}$ (and Hydrogen and Deuterium) is described in two recent publications Bodek and Christy (2022, 2023) (similar fits were also done for other nuclei). The fit includes nuclear elastic form factors, nuclear-excitations, quasielastic scattering, resonance production and inelastic continuum. In the quasielastic region the fit is based on the $\psi^{\prime}$ SuSA (Super Scaling Approximation) formalism Maieron et al. (2002); Amaro et al. (2005, 2020); Megias et al. (2014) with Rosenfelder Rosenfelder (1980) Pauli suppression. For the QE region the fit also includes parametrizations of a $multiplicative$ “Longitudinal Quenching Factor” at low $\bf q$ and an $additive$ “Transverse Enhancement” contribution, which is significant at intermediate $\bf q$. In the inelastic resonance and continuum region the fit is based on a Gaussian Fermi motion smeared free proton and neutron cross sections with a multiplicative medium modification factor. The relative normalizations of different ${\rm{}^{12}C}$ cross section data sets are also extracted from the fit. In summary the fit includes:
@@
    Coulomb corrections Gueye et al. (1999) using the Effective Momentum Approximation (EMA) in modeling scattering from nuclear targets. We use $V_{eff}$= 3.10$\pm$ 0.25 MeV for $\rm{}^{12}C$
@@
    A $\bf q$ dependent $E_{\mathrm{shift}}^{QE}(\bf q)$ parameter for the QE process to account for the optical potential Bodek and Cai (2019, 2020) of final state nucleons. The separation energies of a proton and a neutron from $\rm{}^{12}C$ are 16 MeV and 18.7 MeV, respectively. Therefore, we ensure that the QE cross section on protons is zero for excitation energy below 16 MeV, and the QE cross section on neutrons is zero for excitation energy less than 18.7 MeV.
@@
The use of the Christy-Bodek universal fit Bodek and Christy (2022, 2023) in identifying the few data sets that are inconsistent with all the other measurements is illustrated in the top six sub-figures of Fig. 3. Here, the comparison of the fit to recent CLAS-e4nu:2023 cross sectionIsaacson et al. (2023) measurement (which was not included in the universal fit) at E0=1.159 GeV and $\theta$=37.50 shown in sub-figure (a) indicates that normalization of this data relative to all previous data is 0.85. A comparison to the Sealock:1989 measurement Sealock et al. (1989) at 1.108 GeV and $\theta$=37.50 (which was included in the universal fit) shown in sub-figure (b) indicates a normalization of 1.06. A comparison to a recent Mainz:2024Mihovilovič et al. (2024) measurement at 0.855 GeV and $\theta$=700 (which was not included in the universal fit) shown in sub-figure (c) indicates a normalization of 1.03. Comparisons to Zeller:1973 measurementsZeller (1973a); Heimlich et al. (1974) for incident energies of 2.0, 2.5 and 2.7 GeV at $\theta$=37.50 (without any normalizations) are shown in sub-figures (d-f). The Zeller:1973 measurements are excluded from the universal fit because of inconsistent normalizations between the three energies, and an unexplained shift in $\nu$ in the 2.7 GeV data.
@@
In order to minimize the “bin centering” corrections it is critical to bin in a kinematic variable for which features (i.e. peaks) in the cross section remain fixed independent of the angle. For this reason we perform the analysis in bins of both the excitation energy and bins of the square of the invariant final state mass ($W^{2}=M^{2}+2M\nu-Q^{2}$). We use the results of the analysis in bins of excitation energy for $E_{x}$ less than 50 MeV (a region dominated by nuclear excitations) and the results of the analysis in bins of $W^{2}$ for $E_{x}$ greater than 50 MeV (a region dominated by QE scattering and pion production). Afterwards, we convert the $E_{x}$ and $W^{2}$ values in the center of each bin to the corresponding values of $\nu$ in the center of the bin.
The longitudinal and transverse response functions (defined in section II.3) in the nuclear excitation region are well described by the universal fit Bodek and Christy (2023) to measured transverse and longitudinal form factors for each nuclear excitation. The ${\rm{}^{12}C}$ ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ measurements published in Yamaguchi:71 Yamaguchi et al. (1971) in the nuclear excitation region above the proton separation energy (16$<E_{x}<40$ MeV) are compared to the ${\cal R}_{L}/Z^{2}$ and ${\cal R}_{T}/Z^{2}$ predictions from the fit in Figure 1.
@@
As shown in the bottom right panel of Fig. 3 the contribution from nuclear excitation is significant at low $Q^{2}$ (or $\bf q$). For $E_{x}<$16 MeV we use ${\cal R}_{L}$ and ${\cal R}_{T}$ from overall fits Bodek and Christy (2023) to the nuclear excitation form factors (shown as the red solid line in Figure 1). For $4016<E_{x}<$ 40 MeV (when available) we use the precise (±$\pm 3\%$) Yamaguchi:71 measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$. In addition, for ${\cal R}_{T}$, we also use electron scattering data at 180 Goldemberg and Barber (1964); Deforest et al. (1965) as shown in the right top panel of Fig. 5. For $E_{x}>30$ MeV, we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from our analysis of all available electron scattering data as described below.
@@
Figure 5: Nuclear excitation region: Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region ($4016<E_{x}<$ 40 MeV) for ${\rm{}^{12}C}$. The measurements are from Yamaguchi:71 Yamaguchi et al. (1971) except for values of ${\cal R}_{T}(Q^{2}=0)$ which are extracted from photo-absorption data, and values of ${\cal R}_{T}(Q^{2}=0.01)$ which are extracted from electron scattering cross section measurements at 180 (published by Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)).
@@
12 E04-001-2005:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 0.9981 0.0067
13 E04-001-2007:24 Alsalmi et al. (2024); Albayrak et al. (2024); Alsalmi (2019); Alsalmi et al. (2025) 1.0029 0.0070
@@
25 CLAS-e4nu:2023Isaacson et al. (2023) (not used) 0.85 0.02
@@
Table 2: A summary table of the $\rm{}^{12}C$ data sets used in the Christy-Bodek universal fit and in this analysis. Shown are the normalization factors and uncertainties. Data sets 18–23 ( have larger normalization uncertainty) are early data at very low ${\bf q}$. Data sets which are inconsistent with other data sets are identified as not used.
@@ II Inclusive Electron-Nucleon Scattering
$\nu_{elastic}=E_{0}-\frac{E_{0}}{1+2E_{0}{\sin}^{2}\frac{\theta}{2}/M_{A}},$ (4)
and $M_{A}$ is the mass of the nuclear target. For carbon $M_{A}=12u=11.178$ GeV (1 u = 931.502 MeV). Or equivalently
$E_{x}=\nu-\frac{Q^{2}_{elastic}}{2M_{A}},$ (5)
where $Q^{2}_{elastic}$ is the $Q^{2}$ for elastic scattering from a carbon nucleus for incident energy $E_{0}$ and scattering angle $\theta$.
@@ II.3.1 Extraction of response functions from photo-absorption cross sections
The measured photo-absorption cross section Bezić et al. (1969); Ahrens et al. (1975); Carrasco and Oset (1992); Bianchi et al. (1996); Muccifora et al. (1999) on 12C is shown in the top panel of Fig 6, while ${\cal R}_{T}(Q^{2}=0,\nu)$ extracted via Eq 24 is shown in the bottom panel. In electron scattering $Q^{2}$=${\bf q}^{2}+\nu^{2}$ must be $\geq 0$. Therefore, for a fixed $\bf q$ the maximum $\nu$ ($\nu^{max}$) occurs when $\nu=\bf q$.

### v3

@@ I Introduction
Electron scattering cross sections on nuclear targets are completely described by longitudinal (${\cal R}_{L}$) and transverse (${\cal R}_{T}$) nuclear electromagnetic response functions. Here ${\cal R}_{L}$ and ${\cal R}_{T}$ are functions of the energy transfer $\nu$ (or excitation energy $E_{x}$) and the square of the 4-momentum transfer $Q^{2}$ (or alternatively the 3-momentum transfer $\bf q$). Recent theoretical  Mihaila and Heisenberg (2000); Lovato et al. (2016, 2023a); Cloët et al. (2016); Franco-Munoz et al. (2022, 2023); Sobczyk et al. (2020); Carlson et al. (2015, 2002) calculations of ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ can be tested by comparing the predictions to experimental data. We perform extraction of both response functions as functions of (${\bf q},\nu$), as well as ($Q^{2},\nu)$ from an analysis of all available inclusive cross section data for $\rm{}^{12}C$, and compare them to theoretical predictions.
@@
In this paper we focus on the nuclear excitation, quasielastic (QE) and two nucleon final state (2p2h) regions and compare the extracted $\rm{}^{12}C$ ${\cal R}_{L}({\bf q},\nu)$ and ${\cal R}_{T}({\bf q},\nu)$ to nuclear theory predictions of the “Green’s Function Monte Carlo” (GFMC) Lovato et al. (2016, 2023a), the “Energy Dependent-Relativistic Mean Field” (ED-RMF) Franco-Munoz et al. (2022, 2023), the “Short Time Approximation Quantum Monte Carlo” (STA-QMC) Andreoli et al. (2024); Pastore et al. (2020), the “Correlated Fermi Gas” (CFG) Bhattacharya et al. (2024), the NuWro theoretical approach NuW (2024), the achilles Isaacson et al. (2023) theoretical approach, and an improved superscaling model (SuSAv2)Gonzaléz-Jiménez et al. (2014); Megias et al. (2016); Gonzalez-Rosa et al. (2022, 2023). We have not done comparisons to the gibuu theoretical approach because nuclear response function predictions by gibuu are not readily available.
The theoretical approaches of NuWro and achilles  have been implemented in the corresponding neutrino MC generators. Recently, ED-RMF has been implementedMcKean (2024) in an update of the neut neutrino MC generator. The improved SuSAv2 predictions for the 1p1h and 2p2h channels has been implemented in genie. SuSAv2-inelastic is being implemented in genie, and STA-QMC for $\rm{}^{4}He$ has been implemented in genie. In this paper we focus on testing theoretical predictions. The testing of the implementation of the theoretical approaches into the genie neutrino MC generator (run in electron scattering mode) will be presented in a future communication. Similar investigation of electron scattering data on $\rm{}^{40}Ca$ and $\rm{}^{56}Fe$ (as well as hydrogen and deuterium data) are currently under way.
@@
In this communication we report on the extraction of ${\cal R}_{L}$ and ${\cal R}_{T}$ for $\rm{}^{12}C$} at 18 distinct values of both q and $Q^{2}$ by including all available electron scattering cross section measurements on carbon. We note that a significant number of cross section measurements were not available in tabular form and required digitization of the cross sections from figures in the publications. As summarized in Table 1, the response functions are extracted for a large range of energy transfer $\nu$ at 18 fixed values of $\bf q$ in the range $0.1<{\bf q}<3.75$  GeV and at 18 fixed values of $Q^{2}$ in the range $0<Q^{2}<3.5$ GeV2. The range of $\nu$ spans the nuclear excitation, QE, resonance and inelastic continuum. The complete data set consists of about 10,000 $\rm{}^{12}C$} differential cross section measurements points as well as photo-absorption data ($Q^{2}=0$). We also include high precision cross section measurements from Jefferson Lab Hall C experiment E-04-001 Alsalmi et al. (2025); Alsalmi (2019).
@@
In the nuclear excitation region (excitation energy less than 50 MeV) we extract ${\cal R}_{L}(\bf q,\nu)$ and ${\cal R}_{T}(\bf q,\nu)$ from fits to $\rm{}^{12}C$} nuclear excitation form factors by Bodek and Christy Bodek and Christy (2023). In addition, for $16<E_{x}<40$ MeV, values of ${\cal R}_{L}({\bf q},E_{x})$ and ${\cal R}_{T}({\bf q},E_{x})$ are available from the analysis of Yamaguchi:71 Yamaguchi et al. (1971) for five values of $\bf q$ (0.148, 0.167, 0.205, 0.240 and 0.307 GeV). Comparisons of the fit to the Yamaguchi response functions and cross sections are shown in Figures 1 and 2, respectively.
@@ I.1 Overview of this analysis
The Christy-Bodek 2024 universal fit to all available cross section measurements for $\rm{}^{12}C$} (and Hydrogen and Deuterium) is described in two recent publications Bodek and Christy (2022, 2023) (similar fits were also done for other nuclei). The fit includes nuclear elastic form factors, nuclear-excitations, quasielastic scattering, resonance production and inelastic continuum. In the quasielastic region the fit is based on the $\psi^{\prime}$ SuSA (Super Scaling Approximation) formalism Maieron et al. (2002); Amaro et al. (2005, 2020); Megias et al. (2014) with Rosenfelder Rosenfelder (1980) Pauli suppression. For the QE region the fit also includes parametrizations of a multiplicativemultiplicative “Longitudinal Quenching Factor” at low $\bf q$ and an additiveadditive “Transverse Enhancement” contribution, which is significant at intermediate $\bf q$. In the inelastic resonance and continuum region the fit is based on a Gaussian Fermi motion smeared free proton and neutron cross sections with a multiplicative medium modification factor. The relative normalizations of different $\rm{}^{12}C$} cross section data sets are also extracted from the fit. In summary the fit includes:
@@
[…]
@@
    A $\bf q$ dependent $E_{\mathrm{shift}}^{QE}(\bf q)$ parameter for the QE process to account for the optical potential Bodek and Cai (2019) of final state nucleons. The separation energies of a proton and a neutron from $\rm{}^{12}C$ are 16 MeV and 18.7 MeV, respectively. Therefore, we ensure that the QE cross section on protons is zero for excitation energy below 16 MeV, and the QE cross section on neutrons is zero for excitation energy less than 18.7 MeV.
@@
[…]
  15. 15.
    There is an apparent shift of the peak of the $\Delta$(1232) to smaller values of $\nu$ which in the 2024 version of the fit is accounted for by an “effective optical potential”Bodek and Cai (2020).
  16. 16.
@@
The use of the Christy-Bodek 2024 universal fit Bodek and Christy (2022, 2023) in identifying the few data sets that are inconsistent with all the other measurements is illustrated in the top six sub-figures of Fig. 3. Here, the comparison of the fit to recent CLAS-e4nu:2023 cross sectionIsaacson et al. (2023) measurement (which was not included in the universal fit) at E0=1.159 GeV and $\theta$=37.50 shown in sub-figure (a) indicates that normalization of this data relative to all previous data is 0.85. A comparison to the Sealock:1989 measurement Sealock et al. (1989) at 1.108 GeV and $\theta$=37.50 (which was included in the universal fit) shown in sub-figure (b) indicates a normalization of 1.06. A comparison to a recent Mainz:2024Mihovilovič et al. (2024) measurement at 0.855 GeV and $\theta$=700 (which was not included in the universal fit) shown in sub-figure (c) indicates a normalization of 1.03. Comparisons to Zeller:1973 measurementsZeller (1973a); Heimlich et al. (1974) for incident energies of 2.0, 2.5 and 2.7 GeV at $\theta$=37.50 (without any normalizations) are shown in sub-figures (d-f). The Zeller:1973 measurements are excluded from the universal fit because of inconsistent normalizations between the three energies, and an unexplained shift in $\nu$ in the 2.7 GeV data.
@@
In order to minimize the “bin centering” corrections it is critical to bin in a kinematic variable for which features (i.e. peaks) in the cross section remain fixed independent of the angle. For this reason we perform the analysis in bins of both the excitation energy and bins of the square of the final state mass ($W^{2}=M^{2}+2M\nu-Q^{2}$). We use the results of the analysis in bins of excitation energy for $E_{x}$ less than 50 MeV (a region dominated by nuclear excitations) and the results of the analysis in bins of $W^{2}$ for $E_{x}$ greater than 50 MeV (a region dominated by QE scattering and pion production). Afterwards, we convert the $E_{x}$ and $W^{2}$ values in the center of each bin to the corresponding values of $\nu$ in the center of the bin.
[…]
@@
As shown in the bottom right panel of Fig. 3 the contribution from nuclear excitation is significant at low $Q^{2}$ (or $\bf q$). For $E_{x}<$16 MeV we use ${\cal R}_{L}$ and ${\cal R}_{T}$ from overall fits Bodek and Christy (2023) to the nuclear excitation form factors (shown as the red solid line in Figure 1). For $16<E_{x}<40$ MeV (when available) we use the precise ($\pm 3\%$) Yamaguchi:71 measurements of ${\cal R}_{L}$ and ${\cal R}_{T}$. In addition, for ${\cal R}_{T}$, we also use electron scattering data at 180 Goldemberg and Barber (1964); Deforest et al. (1965) as shown in the right top panel of Fig. 5. For $E_{x}>30$ MeV, we extract ${\cal R}_{L}$ and ${\cal R}_{T}$ from our analysis of all available electron scattering data as described below.
@@
Figure 5: Nuclear excitation region: Extractions of ${\cal R}_{L}$ and ${\cal R}_{T}$ in the nuclear excitation region ($16<E_{x}<40$ MeV) for $\rm{}^{12}C$}. The measurements are from Yamaguchi:71 Yamaguchi et al. (1971) except for values of ${\cal R}_{T}(Q^{2}=0)$ which are extracted from photo-absorption data, and values of ${\cal R}_{T}(Q^{2}=0.01)$ which are extracted from electron scattering cross section measurements at 180 (published by Goldemberg:64 Goldemberg and Barber (1964) and Deforest:65 Deforest et al. (1965)).
@@
12 E04-001-2005:24 Alsalmi (2019); Alsalmi et al. (2025) 0.9981 0.0067
13 E04-001-2007:24 Alsalmi (2019); Alsalmi et al. (2025) 1.0029 0.0070
@@
25 CLAS-e4nu:2023Isaacson et al. (2023) 0.85 0.02
26 Garino::1992Garino et al. (1992) (not used) 1.00 0.02
27 Ricco:1968:Ricco:1968bsq (not used) - -
@@
Ahrens:75 Ahrens et al. (1975)(photoabsorption) - -
(not-included-in-fit)
Carrasco:89 Carrasco and Oset (1992)(photoabsorption) - -
Bianchi:95 Bianchi et al. (1996)(photoabsorption) - -
Bezic:69 Bezić et al. (1969)(photoabsorption) - -
(not-included-in-fit)
Table 2: A summary table of the $\rm{}^{12}C$ data sets used in the Christy-Bodek 2024 universal fit and in this analysis. Shown are the normalization factors and uncertainties. Data sets 18–23 ( have larger normalization uncertainty) are early data at very low $\bf q$}. Data sets which are inconsistent with other data sets are identified as not used.
@@ II Inclusive Electron-Nucleon Scattering
$\nu_{elastic}=E_{0}-\frac{E_{0}}{1+2E_{0}{\sin}^{2}\frac{\theta}{2}/M_{A}}=\frac{Q^{2}_{elastic}}{2M_{A}},$ (4)
where $Q^{2}_{elastic}$ is the $Q^{2}$ for elastic scattering from a carbon nucleus for incident energy $E_{0}$ and scattering angle $\theta$ and $M_{A}$ is the mass of the nuclear target. For carbon $M_{A}=12u=11.178$ GeV (1 u = 931.502 MeV). The excitation energy is
$E_{x}=\nu-\nu_{elastic}$ (5)
Alternatively, one can describe nuclear excitations in terms of the mass of the excited carbon nucleus.
$W^{2}_{A}=M^{2}_{A}+2M_{A}\nu-Q^{2},$ (6)
@@ II.3.1 Extraction of response functions from photo-absorption cross sections
The measured photo-absorption cross section Bezić et al. (1969); Ahrens et al. (1975); Carrasco and Oset (1992); Bianchi et al. (1996); Muccifora et al. (1999) on 12C is shown on the top panel of Fig 6. The dashed red line is the expectation from the Quasi-deuteron model Plujko et al. (2018)(scattering from short-correlation neutron-proton pairs). The Values of ${\cal R}_{T}(Q^{2}=0,\nu)$ extracted via Eq 25 is shown in the bottom panel. In electron scattering $Q^{2}$${\bf q}^{2}+\nu^{2}$ must be $\geq 0$. Therefore, for a fixed $\bf q$ the maximum $\nu$ ($\nu^{max}$) occurs when $\nu=\bf q$ and